# Rational Metareasoning for Large Language     Models

### v2

@@ 7 Conclusion
In this work, we introduced a novel method inspired by previous research in AI and cognitive science aimed at reducing inference costs in LLMs. We operationalize the concept of rational metareasoning – which formalizes humans’ adaptive use of cognitive resources – with a novel reward function based on the Value of Computation (VOC). This reward function trains the LLM to optimize the use of intermediate reasoning steps in task execution, enabling it to balance task performance with computational efficiency. Empirically, we find that this approach significantly reduces the number of generated tokens and input context length while maintaining comparable performance across diverse datasets.
However, we note some limitations to our approach. First, the focus of our work here is enhancing the efficiency of large language models in reasoning tasks, rather than the overall performance of such systems. Thus, we aimed to reduce unnecessary intermediate reasoning steps, rather than improve the quality of reasoning per se. It remains to be seen whether our approach can be extended to enhance task performance in addition to computational efficiency. Additionally, we tested our approach on well-established datasets in science, commonsense reasoning, and math, but it remains untested in more specialized domains. Notably, one of these untested settings is the agentic one, where LLMs act as agents to perform complex tasks in digital environments (Yao et al., 2023b; Schick et al., 2023). Our method could be adapted by including the number of LLM API calls to such tools in the cost section of our reward function, ideally encouraging the model to minimize unnecessary tool calls. Broader testing could help assess its generalizability and effectiveness across diverse contexts. Additionally, the VoC reward function is highly adaptable and can be used with various algorithms. For instance, it can provide numerical rewards for PPO or guide DPO by training on samples with the highest and lowest rewards. In this work, we focused on EI, the most sample-efficient algorithm for reasoning tasks Havrilla et al. (2024), to demonstrate its generalization across models and benchmarks. Exploring its application to other algorithms is an exciting direction for future research.
Most excitingly, our work demonstrates how cognitively-inspired reward functions can endow LLMs with desirable inference-time properties, opening a broad avenue of future work. Given its flexibility, this method could be integrated into instruction tuning to potentially enhance performance, even in scenarios where verifying the correctness of answers is challenging. Since the utility measure within the reward function can be tailored to prioritize any desired, measurable property, this approach offers the potential to guide models toward achieving these enhanced qualities while still benefiting from the reduced computational costs.
Appendix A Full results tables
@@
Phi-2 Direct Few Shot 86.1 ($\pm$ 1.1) 68.4 ($\pm$ 2.6) 10.2 ($\pm$ 1.6) 38.0 ($\pm$ 3.0)
Cot Few Shot 82.8 ($\pm$ 1.2) 66.0 ($\pm$ 2.6) 41.2 ($\pm$ 2.8) 36.3 ($\pm$ 3.1)
STaR 82.9 ($\pm$ 1.1) 63.4 ($\pm$ 2.9) 51.6 ($\pm$ 2.8) 54.8 ($\pm$ 2.8)
Metareasoning 83.9 ($\pm$ 1.3) 61.6 ($\pm$ 2.7) 54.9 ($\pm$ 2.6) 58.2 ($\pm$ 2.9)
Meta-Llama-3-8B Direct Few Shot 86.7 ($\pm$ 1.0) 72.4 ($\pm$ 2.5) 13.1 ($\pm$ 1.8) 49.7 ($\pm$ 3.1)
Cot Few Shot 87.5 ($\pm$ 1.2) 70.6 ($\pm$ 2.5) 44.7 ($\pm$ 2.6) 52.1 ($\pm$ 3.0)
STaR 83.6 ($\pm$ 1.2) 67.4 ($\pm$ 2.7) 39.7 ($\pm$ 2.5) 64.6 ($\pm$ 2.7)
Metareasoning 83.6 ($\pm$ 1.2) 70.5 ($\pm$ 2.7) 40.3 ($\pm$ 2.8) 63.3 ($\pm$ 2.8)
Mistral-7B-v0.3 Direct Few Shot 84.9 ($\pm$ 1.3) 69.6 ($\pm$ 2.6) 10.2 ($\pm$ 1.6) 45.8 ($\pm$ 2.9)
Cot Few Shot 84.3 ($\pm$ 1.2) 66.8 ($\pm$ 2.7) 35.5 ($\pm$ 2.4) 43.3 ($\pm$ 2.9)
STaR 82.9 ($\pm$ 1.4) 69.2 ($\pm$ 2.5) 32.8 ($\pm$ 2.4) 55.4 ($\pm$ 3.0)
Metareasoning 81.3 ($\pm$ 1.2) 69.0 ($\pm$ 2.3) 36.9 ($\pm$ 2.8) 54.9 ($\pm$ 3.2)
@@
Phi-2 Direct Few Shot 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0)
Cot Few Shot 74.6 ($\pm$ 1.4) 80.4 ($\pm$ 2.2) 161.0 ($\pm$ 2.9) 459.6 ($\pm$ 16.5)
STaR 91.4 ($\pm$ 1.0) 80.6 ($\pm$ 1.5) 155.7 ($\pm$ 3.1) 299.9 ($\pm$ 7.2)
Metareasoning 55.2 ($\pm$ 1.2) 53.0 ($\pm$ 1.7) 143.8 ($\pm$ 2.6) 240.4 ($\pm$ 6.8)
@@
Cot Few Shot 59.5 ($\pm$ 0.9) 69.9 ($\pm$ 2.0) 157.1 ($\pm$ 2.8) 307.1 ($\pm$ 11.4)
STaR 61.3 ($\pm$ 1.2) 55.7 ($\pm$ 1.9) 155.3 ($\pm$ 2.8) 206.4 ($\pm$ 6.2)
Metareasoning 31.4 ($\pm$ 0.7) 26.1 ($\pm$ 0.7) 125.8 ($\pm$ 4.7) 199.2 ($\pm$ 8.2)
Mistral-7B-v0.3 Direct Few Shot 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0)
Cot Few Shot 71.0 ($\pm$ 1.4) 66.2 ($\pm$ 2.3) 174.8 ($\pm$ 3.6) 418.8 ($\pm$ 16.9)
STaR 93.1 ($\pm$ 1.3) 85.0 ($\pm$ 1.9) 175.7 ($\pm$ 3.8) 275.9 ($\pm$ 5.7)
Metareasoning 39.7 ($\pm$ 1.0) 25.0 ($\pm$ 0.7) 158.6 ($\pm$ 3.6) 234.3 ($\pm$ 7.2)
@@
Appendix B Alternative plot of accuracy vs cost
Figure 3: Cost and performance. Left panel: the distribution of input and output tokens across various methods and datasets. Our method, Metareasoning, eliminates the need for few-shot prompting (reducing input tokens) and trains the model to use fewer reasoning tokens compared to STaR. Right panel: Metareasoning matches or improves accuracy relative to other methods, despite substantially reducing inference costs. Across datasets, the reduction in token count (left) is inversely related to the effectiveness of the CoT (right): the model learns that CoT is essential in GSM8K and Proofwriter, but less necessary in CommonsenseQA.
Appendix C Additional out of domain testing
To further evaluate the out-of-domain generalization capabilities of our method, we conducted additional assessments on AQUA-RAT (Ling et al., 2017) and MMLU Hendrycks et al. (2021) using the Phi-2 model. For MMLU, we incorporated a few-shot prompting baseline with examples drawn from datasets used during training and compared the results to STaR and Metareasoning. AQUA-RAT was tested under the same conditions. The results indicate that our method performs comparably to few-shot prompting and STaR on both datasets, while requiring fewer output tokens. This analysis highlights the strength of our approach in achieving a balance between computational efficiency and accuracy in OOD scenarios. Comprehensive results for MMLU and AQUA-RAT are provided in Table 6.
Dataset Method Accuracy (%) $\uparrow$ # Input Tokens # Output Tokens
MMLU Direct Few-Shot 34.1 ($\pm$ 1.1) 495.7 ($\pm$ 2.1) 0.0 ($\pm$ 0.0)
CoT Few-Shot 50.7 ($\pm$ 1.3) 1050.7 ($\pm$ 2.2) 99.1 ($\pm$ 2.2)
STaR 48.2 ($\pm$ 1.3) 94.7 ($\pm$ 2.1) 110.5 ($\pm$ 1.5)
Metareasoning 48.3 ($\pm$ 1.4) 94.7 ($\pm$ 2.2) 78.7 ($\pm$ 1.6)
Aqua-Rat Direct Few-Shot 22.4 ($\pm$ 5.1) 477.9 ($\pm$ 2.5) 0.0 ($\pm$ 0.0)
CoT Few-Shot 38.6 ($\pm$ 5.5) 1032.9 ($\pm$ 2.6) 152.6 ($\pm$ 11.3)
STaR 42.9 ($\pm$ 5.9) 76.9 ($\pm$ 2.5) 138.8 ($\pm$ 11.5)
Metareasoning 36.6 ($\pm$ 5.9) 76.9 ($\pm$ 2.4) 111.2 ($\pm$ 7.7)
Table 6: Additional OOD testing on Phi-2. This table compares the performance of our method against STaR, CoT and Direct Few-Shot across two out-of-domain benchmarks: MMLU and AQUA-RAT. The results highlight that our method achieves comparable accuracy while requiring fewer output tokens, demonstrating both efficiency and robustness in diverse reasoning tasks.
Appendix D Comparison to instruction tuned models
It is reasonable to question whether instruction-tuned models can dynamically adjust the length of their reasoning chains based on task complexity. To explore this, we tested Meta-Llama-3-8B-Instruct using the following prompt:
@@ Appendix D Comparison to instruction tuned models
Answer the following question, thinking step by step to get to the answer. You can think however long you need, but answer as soon as you’re ready. Keep you response concise and use the minimum number of steps to get to the answer. Once you’re finished thinking, write your answer after the ’Answer: ’ prompt.
Unfortunately, this method did not effectively adapt response lengths to align with task complexity in our experiments, as shown in Tables 7 and 8 and Fig. 4. While the observed performance drop might partially result from deviations from the expected input format, it is evident that the length of the reasoning chains was generally much higher and not meaningfully adaptive to the complexity of the task.
Model Method Accuracy (%) $\uparrow$ # Input Tokens # Output Tokens
Meta-Llama-3-8B Direct Few-Shot 55.5 ($\pm$ 1.2) 421.3 ($\pm$ 1.0) 1.0 ($\pm$ 0.0)
CoT Few-Shot 63.7 ($\pm$ 1.4) 995.0 ($\pm$ 0.9) 148.4 ($\pm$ 3.0)
CoT Instruct 57.7 ($\pm$ 1.3) 73.3 ($\pm$ 0.9) 190.4 ($\pm$ 2.5)
STaR 63.9 ($\pm$ 1.5) 73.3 ($\pm$ 1.0) 119.7 ($\pm$ 1.9)
Metareasoning 64.4 ($\pm$ 1.4) 73.3 ($\pm$ 0.9) 95.6 ($\pm$ 2.3)
Table 7: Llama Instruct results. The table shows accuracy (in %), average number of input tokens, and average number of output tokens for each method. Despite the instruction, reasoning chains remain lengthy.
Figure 4: Llama Instruct plot. Left panel: the distribution of input and output tokens across various methods and datasets. Right panel: Comparison of performance across datasets.
Model Method Hard Split Easy Split Length Reduction (%) $\uparrow$
Meta-Llama-3-8B CoT Few-Shot 132.0 ($\pm$ 3.0) 57.0 ($\pm$ 1.0) 56.8
CoT Instruct 181.0 ($\pm$ 4.0) 150.0 ($\pm$ 2.0) 17.1
STaR 126.0 ($\pm$ 3.0) 64.0 ($\pm$ 2.0) 49.2
Metareasoning 99.5 ($\pm$ 2.5) 27.0 ($\pm$ 1.0) 72.9
Table 8: The length of generated reasoning chains across models and methods (mean number of tokens with 95% confidence intervals). Adaptive methods should maximize the difference in length between the two distributions. Our method (Metareasoning) reduces the overall length and increases the difference in the length distribution between the Hard and Easy splits (as seen in column Length Reduction), demonstrating an improvement in the model’s ability to adapt the reasoning length to the complexity of the task.
Appendix E Qualitative examples
@@
Below we show one example of question, reasoning chain and answer per dataset, generated with each method: CoT Few Shots (Yellow), STaR (green) and Metareasoning (Red). The following examples were generated by Phi-2 (Javaheripi et al., 2023). The reasoning chains shown here serve as intermediate computations that guide the model to its final answers. While not intended as direct explanations, the examples remain interpretable and demonstrate the effectiveness of the respective methods.
@@
Appendix F Direct Finetuning Baseline
We also tested our method against a direct finetuning baseline. It can be seen that while this method provides an even bigger efficiency gain, the performance is less consistent.
Figure 9:

### v3

@@ 8 Conclusion
We introduced a cognitively inspired reward function grounded in rational metareasoning, aimed at optimizing LLM inference by balancing performance and computational cost. Empirically, we showed that this approach substantially reduces the number of generated tokens and input context length, while consistently improving task accuracy across a diverse range of benchmarks. Most excitingly, our work demonstrates how cognitively-inspired reward functions can endow LLMs with desirable inference-time properties, opening a broad avenue of future work. Given its flexibility, this method could be integrated into instruction tuning to potentially enhance performance, even in scenarios where verifying the correctness of answers is challenging. Since the utility measure within the reward function can be tailored to prioritize any desired, measurable property, this approach offers the potential to guide models toward achieving these enhanced qualities while still benefiting from the reduced computational costs.
Appendix A Complete Results Tables
In this section, we present the comprehensive results tables.
@@ Appendix A Complete Results Tables
Llama-3.2-3B Direct Few Shot 70.1 ($\pm$ 1.8) 56.2 ($\pm$ 3.2) 7.6 ($\pm$ 1.6) 0.0 ($\pm$ 1.3)
Cot Few Shot 63.8 ($\pm$ 1.9) 43.3 ($\pm$ 3.3) 27.7 ($\pm$ 2.4) 4.2 ($\pm$ 4.5)
Direct SFT 82.8 ($\pm$ 1.2) 66.0 ($\pm$ 2.6) 41.2 ($\pm$ 2.8) 36.3 ($\pm$ 3.1)
STaR 66.6 ($\pm$ 1.9) 52.5 ($\pm$ 3.5) 26.8 ($\pm$ 2.5) 43.6 ($\pm$ 4.5)
RaM 71.2 ($\pm$ 1.7) 55.8 ($\pm$ 3.3) 32.3 ($\pm$ 2.5) 41.3 ($\pm$ 4.4)
Llama-3.2-3B-Instruct CoT Prompting 75.9 ($\pm$ 1.7) 58.1 ($\pm$ 3.2) 71.6 ($\pm$ 2.2) 7.3 ($\pm$ 4.4)
Llama-3.1-8B Direct Few Shot 82.6 ($\pm$ 1.5) 66.4 ($\pm$ 3.0) 14.9 ($\pm$ 2.0) 19.4 ($\pm$ 4.4)
Cot Few Shot 80.9 ($\pm$ 1.5) 60.4 ($\pm$ 3.2) 52.8 ($\pm$ 2.6) 32.9 ($\pm$ 4.2)
STaR 82.0 ($\pm$ 1.6) 64.0 ($\pm$ 3.0) 45.9 ($\pm$ 2.8) 55.8 ($\pm$ 4.1)
Metareasoning 80.3 ($\pm$ 1.5) 67.2 ($\pm$ 3.1) 53.7 ($\pm$ 2.5) 52.7 ($\pm$ 4.0)
Llama-3.1-8B-Instruct CoT Prompting 85.4 ($\pm$ 1.4) 66.6 ($\pm$ 3.0) 80.5 ($\pm$ 2.0) 11.2 ($\pm$ 4.3)
@@
Llama-3.2-3B Direct Few Shot 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0)
Cot Few Shot 169.6 ($\pm$ 2.3) 182.0 ($\pm$ 9.7) 236.9 ($\pm$ 8.0) 459.2 ($\pm$ 14.1)
STaR 121.0 ($\pm$ 2.2) 130.0 ($\pm$ 1.4) 259.7 ($\pm$ 9.0) 380.0 ($\pm$ 7.2)
RaM 77.4 ($\pm$ 1.5) 91.3 ($\pm$ 2.6) 184.1 ($\pm$ 7.8) 330.4 ($\pm$ 7.4)
Llama-3.2-3B-Instruct CoT Prompting 230.8 ($\pm$ 2.2) 189.1 ($\pm$ 3.7) 192.1 ($\pm$ 4.6) 424.6 ($\pm$ 11.0)
@@
Cot Few Shot 147.9 ($\pm$ 1.5) 153.6 ($\pm$ 6.8) 232.7 ($\pm$ 8.6) 472.1 ($\pm$ 16.8)
STaR 119.6 ($\pm$ 2.0) 143.6 ($\pm$ 2.3) 256.9 ($\pm$ 7.5) 340.4 ($\pm$ 9.6)
Metareasoning 91.8 ($\pm$ 1.4) 93.1 ($\pm$ 3.2) 171.4 ($\pm$ 5.4) 228.7 ($\pm$ 10.1)
Llama-3.1-8B-Instruct CoT Prompting 233.0 ($\pm$ 2.2) 200.7 ($\pm$ 4.3) 210.8 ($\pm$ 5.0) 435.1 ($\pm$ 11.7)
@@
Appendix B CoT Prompting
For the “Instruct” versions of our models we use the following prompt:
@@ Appendix B CoT Prompting
Answer the following question, thinking step by step to get to the answer. You can think however long you need, but answer as soon as you’re ready. Use the minimum number of steps to get to the answer. Once you’re finished thinking, you must end your response with ’The answer is [X]’, where [X] is the final answer to the question.
Appendix C Qualitative examples
@@
Below we show one example of question, reasoning chain and answer per dataset, generated with each method: CoT Few Shots (blue), STaR (yellow) and RaM (green). The following examples were generated by Llama-3.2-3B (Dubey et al., 2024). The reasoning chains shown here serve as intermediate computations that guide the model to its final answers. While not intended as direct explanations, the examples remain interpretable and demonstrate the effectiveness of the respective methods.